# QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow

### v1

@@ E.1 Reviewing quantum Teichmuller theory and Teichmuller TQFT
In [310], it is shown that the quantum $6j$ symbols in $U_{q}(SL(2,\mathbb{R}))$ are related to the tetrahedra operator $T$ introduced by Kashaev in quantum Teichmuller theory [178, *Kashaev:2000ku, 64].
To appreciate the relation given in [310], we provide a brief review of quantum Teichmuller theory and the $T$ operator introduced in [178, *Kashaev:2000ku]. To ensure convergence within the discrete formulation of Teichmuller TQFT, these $T$ operators were later extended to their charged counterparts, as developed in [24].
@@
The Teichmuller space can be identified with the space of deformations of these constant-curvature metrics. The idea of Penner coordinates for the Teichmuller space is to use geodesic lengths to parameterize the space of metrics[311, *10.4310/jdg/1214448257]. Consider, thus, this specific hyperbolic conformal frame. The punctures are mapped to cusps. In the following, we would be using these two terms interchangeably.
@@
The main idea of such a construction by Kashaev in [178, *Kashaev:2000ku] follows from expressing the Fock variables in terms of canonical variables $p,q$ that satisfy the usual commutation relation $[q,p]=1$.

### v2

@@ E.1 Reviewing quantum Teichmuller theory and Teichmuller TQFT
In [319], it is shown that the quantum $6j$ symbols in $U_{q}(SL(2,\mathbb{R}))$ are related to the tetrahedra operator $T$ introduced by Kashaev in quantum Teichmuller theory [174, *Kashaev:2000ku, 58].
To appreciate the relation given in [319], we provide a brief review of quantum Teichmuller theory and the $T$ operator introduced in [174, *Kashaev:2000ku]. To ensure convergence within the discrete formulation of Teichmuller TQFT, these $T$ operators were later extended to their charged counterparts, as developed in [16].
@@
The Teichmuller space can be identified with the space of deformations of these constant-curvature metrics. The idea of Penner coordinates for the Teichmuller space is to use geodesic lengths to parameterize the space of metrics[320, *10.4310/jdg/1214448257]. Consider, thus, this specific hyperbolic conformal frame. The punctures are mapped to cusps. In the following, we would be using these two terms interchangeably.
@@
The main idea of such a construction by Kashaev in [174, *Kashaev:2000ku] follows from expressing the Fock variables in terms of canonical variables $p,q$ that satisfy the usual commutation relation $[q,p]=1$.